OpenStudy (anonymous):

A recipe from Sally's Chocoholic's cookbook requires 50 6/7 (Mixed Number) grams of chocolate for each chocolate fudge cake. How many greams of chocolate would she need if she made one choco cake for each of her 14 closest friends?

OpenStudy (anonymous):

umm...does she need 14 cakes for each person, or just one cake?

OpenStudy (anonymous):

I'm confused on this question

OpenStudy (anonymous):

1 Cake for each person, so 14 cakes

OpenStudy (anonymous):

dang...they hungry friends

OpenStudy (anonymous):

Haha yeah, this question is on a practice test that i just can't seem to figure out!

OpenStudy (anonymous):

well let's do\[14 * 50 \frac{ 6 }{ 7 }\]

OpenStudy (anonymous):

Wouldn't you have to turn the mixed number into a improper fraction?

OpenStudy (anonymous):

712 grams

OpenStudy (anonymous):

i figured you would do that during your work x]

OpenStudy (anonymous):

UCL how did you get that answer!

OpenStudy (anonymous):

this is why the rules say don't just give the answer -.-

OpenStudy (anonymous):

356/7 is the improper fraction

OpenStudy (anonymous):

\[50 \frac{ 6 }{ 7 }\] \[\frac{ 356 }{ 7 } * \frac{ 14 }{ 1 } = \frac{ 4984 }{ 7 }\] simplify that to 712

OpenStudy (anonymous):

Oh.. that might be what i did wrong, i reduced the 7 to 1 and turned the 14 into 7

OpenStudy (anonymous):

well you can cross cancel

OpenStudy (anonymous):

356 * 2 = 712

OpenStudy (anonymous):

OHHHHH ok, i see where i made my mistake... thank you very much!